# Modeling nonlinear scales with COLA: preparing for LSST-Y1

### v1

@@ Abstract
Year 1 results of the Legacy Survey of Space and Time (LSST) will provide tighter constraints on small-scale cosmology, beyond the validity of linear perturbation theory. This heightens the demand for a computationally affordable prescription that can accurately capture nonlinearities in beyond-$\Lambda$CDM models. The COmoving Lagrangian Acceleration (COLA) method, a cost-effective N-body technique, has been proposed as a viable alternative to high-resolution N-body simulations for training emulators of the nonlinear matter power spectrum. In this study, we evaluate this approach by employing COLA emulators to conduct a cosmic shear analysis with LSST-Y1 simulated data across three different nonlinear scale cuts. We use the $w$CDM model, for which the EuclidEmulator2 (ee2) exists as a benchmark, having been trained with high-resolution N-body simulations. We primarily utilize COLA simulations with mass resolution $M_{\rm part}\approx 8\times 10^{10}~{}h^{-1}M_{\odot}$ and force resolution $\ell_{\rm force}=0.5~{}h^{-1}$pc, though we also test refined settings with $M_{\rm part}\approx 1\times 10^{10}~{}h^{-1}M_{\odot}$ and force resolution $\ell_{\rm force}=0.17~{}h^{-1}$pc. We find the performance of the COLA emulators is sensitive to the placement of high-resolution N-body reference samples inside the prior, which only ensure agreement in their local vicinity. However, the COLA emulators pass stringent criteria in goodness-of-fit and parameter bias throughout the prior, when $\Lambda$CDM predictions of ee2 are computed alongside every COLA emulator prediction, suggesting a promising approach for extended models.
@@ I Introduction
In a weak-lensing analysis, improving small-scale modeling primarily requires refining the prediction of the matter power spectrum, $P(k,z)$. At linear order, this quantity can be quickly evaluated with accuracy using Einstein-Boltzmann codes, such as camb111github.com/cmbant/CAMB and class222github.com/lesgourg/class_public [22] for General Relativity (GR), as well as eftcamb [23] and hiclass [24] for modified gravity theories [25]. However, the same is not true on nonlinear scales, where the equations of motion for the matter density field become coupled, and the linear theory approach of evolving modes independently can no longer be applied. Recent efforts have been put forth to efficiently perform convolution integrals appearing in higher order perturbation theory [26], such as the fastpt code [27, 28], the Effective Field Theory (EFT) approach [29, 30, 31, 32, 33, 34, 35, 36] and its implementation [37, 38, 39], as well as its generalizations for some beyond-$\Lambda$CDM models [40, 41, 42, 43, 39, 44, 45, 46]. While these perturbative templates are becoming computationally less costly, they cover a more limited range of the scales being probed by photometric surveys, having only been validated up to the order of $k\sim 0.2\,h$Mpc-1 at $z=0$ [47, 48, 49, 50, 26].
@@
Arguably, the most accurate way of computing the matter power spectrum far into the nonlinear regime is through the use of high-resolution N-body simulations, which are known to be time-consuming and computationally expensive [^3] An estimate of the time taken and the computational cost of running a high-resolution simulation with stage-IV required specifications is discussed in Section 2.1 of [51].. This expense only increases when running these codes for beyond-$\Lambda$CDM models, such as in modified gravity theories, where the equations of motion for the extra degrees of freedom [52] must be solved in addition to the Poisson equation. At the same time, Markov Chain Monte Carlo (MCMC) methods often require $\mathcal{O}(10^{6})$ computations of the matter power spectrum throughout the parameter space in order to obtain robust constraints in standard weak-lensing analyses, necessitating predictions of the power spectrum in $\mathcal{O}(1)$ seconds. Currently, the halo model provides analytic formulas that can be fit to the results of N-body simulations, allowing for fast predictions of the power spectrum up to scales of $k\sim 10\,h$Mpc-1, such as halofit [53, 54, 55]  [^4] While this work was in its final stages, an improved version of halofit was published [56]., hmcode [57, 58, 59] and the halo reaction model approach react [60, 61].
@@
Alternatively, emulation methods have garnered attention in the literature as a means of quickly replicating the results of N-body simulations, both in $\Lambda$CDM [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81], and in beyond-$\Lambda$CDM models [82, 83, 84, 85, 86, 87, 88, 89, 90]. In particular, the EuclidEmulator2 (ee2[91] and the bacco emulator [92] cover a wide range of parameter space, redshifts, and scales. The former was built with a training set of $127$ high-resolution N-body simulations, run with the GPU accelerated pkdgrav3 code [93], while the latter utilizes the rescaling methodology developed by [94] to predict the nonlinear structure formation of $800$ training cosmologies from 6 high-resolution simulations. Notably, ee2 can reproduce the results of the Euclid training simulations to within $\sim 1\%$ error for $z\leq 3$ and $k\leq 1h\mathrm{Mpc}^{-1}$, effectively serving as a near substitute for a high-resolution N-body simulation in $\Lambda$CDM and minor extensions.
@@
In this paper, we assess the performance of COLA emulators in constraining extended models by conducting LSST-Y1 simulated cosmic shear analyses. We consider the $w$CDM model, where the dark energy equation of state $w$ is a free parameter, as ee2 is trained for this model and provides a suitable benchmark. While the original proposal [95] showed sub-percent agreement between COLA and ee2 in their nonlinear predictions down to scales of $k\sim 1~{}h$Mpc-1 for some selected cosmologies, in the present work we perform stress tests of COLA’s performance under large shifts throughout the $w$CDM parameter space. As beyond-$\Lambda$CDM models introduce extra parameters that will further have their own degeneracies with the $\Lambda$CDM parameters, it is ideal to have robust agreement even in extreme regions of the prior. One of the goals of this analysis is to determine the necessary scale cuts on LSST-Y1 cosmic shear data, such that the COLA-based constraints are equivalent to those obtained with high-resolution N-body methods. We aim to inform future analyses of beyond-$\Lambda$CDM models, where high-resolution simulations are scarce. We also test a variety of strategies for enhancing the performance of the COLA emulators, including increasing the COLA resolution settings, and calibrating the simulations with multiple high-resolution reference samples.
@@ II.1.1 COLA Simulation Settings
The implementation of relativistic corrections at large scales using the N-body gauge approach has also been followed by the Euclid collaboration in the Euclid flagship simulations, as well as in the high-resolution N-body simulations used to construct ee2. The implementation in our COLA simulations follows the procedure outlined in [113]. Our COLA simulations themselves were run using the publicly available cola-fml [^5] https://github.com/HAWinther/FML code, and the power spectra from these simulations were evaluated using the on-the-fly code ComputePowerSpectra available in the cola-fml library.
Due to the finite volume of cosmological N-body simulations, the lowest wavenumbers of the power spectra computed from the particle position are undersampled, and the effects of sample variance become more pronounced. To suppress this effect, we follow the technique proposed in [114, 115], known as “pair-fixing”, which consists of averaging the power spectrum between a pair of simulations with fixed amplitudes of their initial random field realizations, but opposite phases. This strategy has also been followed in the ee2 training simulations.
@@ II.2.1 The Boost Factor: computation and validation
Refer to captionRefer to caption
Figure 1: Bounds on the fractional differences between the boost factors from our COLA training simulations with those of ee2, for each wavenumber at $z=0$, for the $50\%$, $90\%$, and $100\%$ of lowest error cosmologies inside the ee2 parameter limits. In the left panels, we use Eq. 4 to compute the boost from our COLA simulations employing the ee2 reference as the single reference cosmology ($N_{\rm refs}=1$). These results include both COLA precision settings defined in Table 1, and both DE models. The right panels show the effect of using the more general Eq. 5 to calibrate the $w$CDM simulations using additional reference anchors. Considering the expense of high-resolution anchors in extended models, the top right panel shows the $w$CDM default-precision disagreement using only 25 additional references scattered throughout the $w$CDM parameter space. Alternatively, the third and fourth panels on the right display errors when the anchors were confined to the $\Lambda$CDM subspace, where we test 100 additional anchors as existing high-resolution $\Lambda$CDM emulators allow for a much greater number. The second panel on the right indicates that only a slight improvement is gained by spreading the 100 additional anchors throughout the entire $w$CDM volume compared to restricting them to $\Lambda$CDM. For computationally effective ways to increase $N_{\rm refs}^{\Lambda\rm CDM}$, see Sec. III.5 and Appendix B.
@@ II.2.2 Emulating the Boost Factor
Here the superscript denotes that a "smearing" of the BAO oscillations has not been applied, which differs from the approach of the bacco emulator [92]. The BAO smearing procedure used by bacco dampens oscillations in the boost, reducing the features in the quantity to be emulated. We implement this approach by using as an emulation variable the following quantity:
@@
In Fig. 3 we display the fractional difference in the $w$CDM validation set at $z=0$ between the emulated boost, obtained by dividing Eqs. 15-16 by $P_{\rm L}(k,z|\bm{\theta})$, and those computed directly from the simulations using Eq. 5. For the default-precision simulations, we show our emulation errors which use $Q^{\rm S}(k,z)$ as the emulation variable in the middle panel, but also include for visual comparison in the left panel the analogous errors using $Q^{\rm NS}(k,z)$. Comparing the range $0.1h\text{Mpc}^{-1}<k<0.5h\text{Mpc}^{-1}$ between the middle and left panels shows the effect of the smearing algorithm on emulation errors. The GP, NN, and PCE default-precision emulators all perform similarly with emulation errors below $\sim 0.5\%$ up to the chosen cutoff for default-precision, $k_{\rm max}=\pi\,h$Mpc-1777For default-precision emulators, we applied $k_{\rm max}=\pi\,h$Mpc-1 for $z<2$ and $k_{\rm max}=\frac{\pi}{2}\,h$Mpc-1 for $z>2$. Conversely, for the enhanced-precision emulators $k_{\rm max}=2\pi$ $h$Mpc-1 for $z<2$ and $k_{\rm max}=\pi$ $h$Mpc-1 for $z>2$. Shot noise can dominate the power spectrum output from the simulations on small scales. Because power dissipates on larger scales for higher redshifts, we adjust $k_{\rm max}$ for $z>2$, to prevent shot noise subtraction from producing spurious negative power spectrum values at these redshifts..
@@ II.3.1 Simulating Cosmic Shear Data
Our analysis is comprised of a simulated cosmic shear survey based on LSST-Y1, as in [119]. The fiducial data vectors and covariance matrix used in the analysis are computed following the methodology introduced in [120]. The specific survey specifications, such as footprint area and effective number of galaxies per solid angle $n_{\mathrm{eff}}$ are based on the LSST DESC Science Requirement Document [121], and are shown in Table 4. The source galaxy redshifts are drawn from a Smail distribution, $n(z)\propto z^{2}\exp[-(z/z_{0})^{\alpha}]$, normalized by $n_{\mathrm{eff}}$, where $(z_{0},\alpha)=(0.191,0.870)$. The drawn samples are divided into 5 tomographic redshift bins with equal number of galaxies. The bins are then convolved with a Gaussian redshift uncertainty of $0.02(1+z)$. The cosmic shear angular two-point correlation function is calculated from theory in the following way. First, the lensing efficiency $q_{\kappa}^{i}$ is defined as:
@@
  • Intrinsic Alignment: tidal gravitational fields can account for correlations between galaxy shapes that are not caused by weak lensing. To model these correlations, we use the "nonlinear linear alignment" (NLA) model, characterized by an amplitude $a$ and exponent $\eta$ (see, e.g. [124, 125]). In this model, the intrinsic shape field $\gamma_{ij}^{\mathrm{IA}}$ is given by:
    $\gamma_{ij}^{\mathrm{IA}}=-a_{1}\bar{C}_{1}\frac{\rho_{\mathrm{c}}\Omega_{m}}{D(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta_{1}},$ (26)
    where $\bar{C}_{1}=5\times 10^{-14}\mathrm{M}_{\odot}h^{-2}\mathrm{Mpc}^{2}$ is a normalization constant, $\rho_{\mathrm{c}}=3H_{0}^{2}/8\pi G$ is the critical density, $D(z)$ is the growth factor, $z_{0}=0.62$ is a pivot redshift and $a_{1}$ and $\eta_{1}$ are model parameters.

### v2

@@ Abstract
Year 1 results of the Legacy Survey of Space and Time (LSST) will provide tighter constraints on small-scale cosmology, beyond the validity of linear perturbation theory. This heightens the demand for a computationally affordable prescription that can accurately capture nonlinearities in beyond-$\Lambda$CDM models. The COmoving Lagrangian Acceleration (COLA) method, a cost-effective N-body technique, has been proposed as a viable alternative to high-resolution N-body simulations for training emulators of the nonlinear matter power spectrum. In this study, we evaluate this approach by employing COLA emulators to conduct a cosmic shear analysis with LSST-Y1 simulated data across three different nonlinear scale cuts. We use the $w$CDM model, for which the EuclidEmulator2 (ee2) exists as a benchmark, having been trained with high-resolution N-body simulations. We primarily utilize COLA simulations with mass resolution $M_{\rm part}\approx 8\times 10^{10}~{}h^{-1}M_{\odot}$ and force resolution $\ell_{\rm force}=0.5~{}h^{-1}$pc, though we also test refined settings with $M_{\rm part}\approx 1\times 10^{10}~{}h^{-1}M_{\odot}$ and force resolution $\ell_{\rm force}=0.17~{}h^{-1}$pc. We find the performance of the COLA emulators is sensitive to the placement of high-resolution N-body reference samples inside the prior, which only ensure agreement in their local vicinity. However, the COLA emulators pass stringent criteria in goodness-of-fit and parameter bias throughout the prior, when existing high-resolution $\Lambda$CDM emulators are leveraged alongside the COLA emulators to predict the respective $\Lambda$CDM parameters, suggesting a promising template for extensions to $\Lambda$CDM.
@@ I Introduction
In a weak-lensing analysis, improving small-scale modeling primarily requires refining the prediction of the matter power spectrum, $P(k,z)$. At linear order, this quantity can be quickly evaluated with accuracy using Einstein-Boltzmann codes, such as camb111github.com/cmbant/CAMB and class222github.com/lesgourg/class_public [22] for General Relativity (GR), as well as eftcamb [23] and hiclass [24] for modified gravity theories [25]. However, the same is not true on small scales, where the linear theory assumption of small matter overdensities breaks down, so that alternative methods are required to accurately compute the matter power spectrum.
Recent efforts have been put forth to efficiently perform convolution integrals appearing in higher order perturbation theory [26], such as the Effective Field Theory (EFT) approach [27, 28, 29, 30, 31, 32, 33, 34] and its implementation [35, 36, 37, 38, 39], as well as its generalizations for some beyond-$\Lambda$CDM models [40, 41, 42, 43, 37, 44, 45, 46]. While these perturbative templates are becoming computationally less costly, they cover a more limited range of scales, having been validated up to the order of $k\sim 0.2\,h$Mpc-1 at $z=0$ [47, 48, 49, 50, 26]. However, for photometric probes, the weak lensing kernel is sensitive to sufficiently smaller scales, limiting the applicability of these approaches.
@@
Arguably, the most accurate way of computing the matter power spectrum far into the nonlinear regime is through the use of high-resolution N-body simulations, which are known to be time-consuming and computationally expensive [^3] An estimate of the time taken and the computational cost of running a high-resolution simulation with stage-IV required specifications is discussed in Sec. 2.1 of [51].. This expense only increases when running these codes for beyond-$\Lambda$CDM models, such as in modified gravity theories, where the equations of motion for the extra degrees of freedom [52] must be solved in addition to the Poisson equation. At the same time, Markov Chain Monte Carlo (MCMC) methods often require $\mathcal{O}(10^{6})$ computations of the matter power spectrum throughout the parameter space in order to obtain robust constraints in standard weak-lensing analyses, necessitating predictions of the power spectrum in $\mathcal{O}(1)$ seconds. Currently, the halo model provides analytic formulas that can be fit to the results of N-body simulations, allowing for fast predictions of the power spectrum up to scales of $k\sim 10\,h$Mpc-1, such as halofit [53, 54, 55]  [^4] While this work was in its final stages, an improved version of halofit was published [56]., hmcode [57, 58, 59] and the halo reaction model approach react [60, 61].
@@
Alternatively, emulation methods have garnered attention in the literature as a means of quickly replicating the results of N-body simulations, both in $\Lambda$CDM [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81], and in beyond-$\Lambda$CDM models [82, 83, 84, 85, 86, 87, 88, 89, 90]. In particular, the EuclidEmulator2 (ee2[91], the bacco emulator [92] and the aemulus emulator [93] cover a wide range of parameter space, redshifts, and scales. ee2 was built with a training set of $127$ high-resolution N-body simulations, run with the GPU accelerated pkdgrav3 code [94]. The bacco emulator, utilizes the rescaling methodology developed by [95] to predict the nonlinear structure formation of $800$ training cosmologies from 6 high-resolution simulations, while aemulus was trained with 150 simulations using the gadget-3 code. Such emulators serve as a near substitute for a high-resolution N-body simulation in $\Lambda$CDM and minor extensions, for example ee2 can reproduce the results of the Euclid training simulations to within $\sim 1\%$ error for $z\leq 3$ and $k\leq 1h\mathrm{Mpc}^{-1}$.
@@
In this paper, we assess the performance of COLA emulators in constraining extended models by conducting LSST-Y1 simulated cosmic shear analyses. We consider the $w$CDM model, where the dark energy equation of state $w$ is a free parameter, as emulators trained on high-resolution simulations exist for this model and provide a suitable benchmark. As there are known disagreements between existing emulators in the literature on small-scales [^5] See for instance Sec. 8 of [98], we choose ee2 as the single baseline with which we measure the performance of the COLA emulators. Consequently, our methodology is tailored to target agreement between the COLA emulators and ee2 specifically, however we establish a general template within which one can in principle substitute a preferred emulator or other nonlinear prescription to target.
While the original proposal [96] showed sub-percent agreement between COLA and ee2 in their nonlinear predictions down to scales of $k\sim 1~{}h$Mpc-1 for some selected cosmologies, in the present work we perform stress tests of COLA’s performance under large shifts throughout the $w$CDM parameter space. As beyond-$\Lambda$CDM models introduce extra parameters that will further have their own degeneracies with the $\Lambda$CDM parameters, it is ideal to have robust agreement even in extreme regions of the prior.
One of the goals of this analysis is to determine the necessary scale cuts on LSST-Y1 cosmic shear data, such that the COLA-based constraints are equivalent to those obtained with high-resolution N-body methods. We aim to inform future analyses of beyond-$\Lambda$CDM models, where high-resolution simulations are scarce. We also test a variety of strategies for enhancing the performance of the COLA emulators, including increasing the COLA resolution settings, and calibrating the simulations with multiple high-resolution reference samples.
@@ II.1.1 COLA Simulation Settings
The implementation of relativistic corrections at large scales using the N-body gauge approach has also been followed by the Euclid collaboration in the Euclid flagship simulations, as well as in the high-resolution N-body simulations used to construct ee2. The implementation in our COLA simulations follows the procedure outlined in [110, 111]. Our COLA simulations themselves were run using the publicly available cola-fml [^6] https://github.com/HAWinther/FML code, and the power spectra from these simulations were evaluated using the on-the-fly code ComputePowerSpectra available in the cola-fml library.
Due to the finite volume of cosmological N-body simulations, the lowest wavenumbers of the power spectra computed from the particle position are undersampled, and the effects of sample variance become more pronounced. There are a number of methods used to suppress the variance effect observed. As an example, we have the technique proposed in [115, 116], known as “pair-fixing”, which consists of averaging the power spectrum between a pair of simulations with fixed amplitudes of their initial random field realizations, but opposite phases. This strategy has also been followed in the ee2 training simulations.
Other interesting avenues to suppress cosmic variance exist, such as the control variates technique [117], which consists of reducing the variance of a random variable by considering another random variable with known mean that is highly correlated with the former. The first usage of this technique in the context of cosmology was in [118, 119, 120], where the authors considered COLA and fastpm simulations as the correlated random variables with known mean to reduce the variance from high resolution N-body simulations. Later on, the authors in [121, 122] showed that one can safely use the Zel’dovich approximation as the control variate to reduce the variance of simulations. However, in this work we have chosen to adopt pair-fixing in order to follow the strategy of our benchmark, ee2.
@@ II.2.1 The Boost Factor: computation and validation
Figure 1: Bounds on the fractional differences between the boost factors from our COLA training simulations with those of ee2, for each wavenumber at $z=0$, for the $50\%$, $90\%$, and $100\%$ of lowest error cosmologies inside the ee2 parameter limits. In the left panels, we use Eq. 4 to compute the boost from our COLA simulations employing the ee2 reference as the single reference cosmology ($N_{\rm refs}=1$). These results include both COLA precision settings defined in Table 1, and both DE models. The right panels show the effect of using the more general Eq. 5 to calibrate the $w$CDM simulations using additional reference anchors. Considering the expense of high-resolution anchors in extended models, the top right panel shows the $w$CDM default-precision disagreement using only 25 additional references scattered throughout the $w$CDM parameter space. Alternatively, the third and fourth panels on the right display errors when the anchors were confined to the $\Lambda$CDM subspace, where we test 100 additional anchors as existing high-resolution $\Lambda$CDM emulators allow for a much greater number. The second panel on the right indicates that only a slight improvement is gained by spreading the 100 additional anchors throughout the entire $w$CDM volume compared to restricting them to $\Lambda$CDM. For similar comparisons between our COLA boosts and ee2 at higher redshifts see Appendix C.
For computationally effective ways to increase $N_{\rm refs}^{\Lambda\rm CDM}$, see Sec. III.5 and Appendix B.
@@ II.2.2 Emulating the Boost Factor
Here the superscript denotes that a "smearing" of the BAO oscillations has not been applied, which differs from the approach of the bacco emulator [92]. The BAO smearing procedure used by bacco dampens oscillations in the boost which are due to large scale bulk flow differences around the BAO peaks between linear and nonlinear prescriptions. This smearing reduces the features in the quantity to be emulated. In our approach, much of this mismatch of the BAO is already mitigated by using Eq. 4. Nevertheless, we implement this procedure by using as an emulation variable the following quantity:
@@
In Fig. 3 we display the fractional difference in the $w$CDM validation set at $z=0$ between the emulated boost, obtained by dividing Eqs. 15-16 by $P_{\rm L}(k,z|\bm{\theta})$, and those computed directly from the simulations using Eq. 5. For the default-precision simulations, we show our emulation errors which use $Q^{\rm S}(k,z)$ as the emulation variable in the middle panel, but also include for visual comparison in the left panel the analogous errors using $Q^{\rm NS}(k,z)$. Comparing the range $0.1h\text{Mpc}^{-1}<k<0.5h\text{Mpc}^{-1}$ between the middle and left panels shows the modest effect of the smearing algorithm on emulation errors. Therefore, we stress that the smearing procedure is not a paramount step that needs to be performed when following our emulation strategy. The GP, NN, and PCE default-precision emulators all perform similarly with emulation errors below $\sim 0.5\%$ up to the chosen cutoff for default-precision, $k_{\rm max}=\pi\,h$Mpc-1888For default-precision emulators, we applied $k_{\rm max}=\pi\,h$Mpc-1 for $z<2$ and $k_{\rm max}=\frac{\pi}{2}\,h$Mpc-1 for $z>2$. Conversely, for the enhanced-precision emulators $k_{\rm max}=2\pi$ $h$Mpc-1 for $z<2$ and $k_{\rm max}=\pi$ $h$Mpc-1 for $z>2$. Shot noise can dominate the power spectrum output from the simulations on small scales. Because power dissipates on larger scales for higher redshifts, we adjust $k_{\rm max}$ for $z>2$, to prevent shot noise subtraction from producing spurious negative power spectrum values at these redshifts..
@@ II.3.1 Simulating Cosmic Shear Data
The weak gravitational lensing of light in the large-scale structure of the Universe distorts galaxies’ shapes, imposing measurable correlations that depend on the matter power spectrum (see [125] for a detailed review). Our cosmic shear analysis is comprised of a simulated survey based on LSST-Y1, as in [126]. The fiducial data vectors and covariance matrix used in the analysis are computed following the methodology introduced in [127]. The specific survey specifications, such as footprint area and effective number of galaxies per solid angle $n_{\mathrm{eff}}$ are based on the LSST DESC Science Requirement Document [128], and are shown in Table 4. The source galaxy redshifts are drawn from a Smail distribution, $n(z)\propto z^{2}\exp[-(z/z_{0})^{\alpha}]$, normalized by $n_{\mathrm{eff}}$, where $(z_{0},\alpha)=(0.191,0.870)$. The drawn samples are divided into 5 tomographic redshift bins with equal number of galaxies. The bins are then convolved with a Gaussian redshift uncertainty of $0.02(1+z)$. The cosmic shear angular two-point correlation function is calculated from theory in the following way. First, the lensing efficiency $q_{\kappa}^{i}$ is defined as:
@@
  • Intrinsic Alignment (IA): tidal gravitational fields can align galaxy shapes, accounting for correlations that are not caused by weak lensing. To model these correlations, we use the "nonlinear linear alignment" (NLA) model, where we use a linear IA model sourced by the nonlinear matter power spectrum [^9] The same nonlinear power spectrum used to calculate the data vector.. The model characterized by an amplitude $a$ and exponent $\eta$ (see, e.g. [131, 132]). In this model, the intrinsic shape field $\gamma_{ij}^{\mathrm{IA}}$ is given by:
    $\gamma_{ij}^{\mathrm{IA}}=-a_{1}\bar{C}_{1}\frac{\rho_{\mathrm{c}}\Omega_{m}}{D(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta_{1}}\left(k_{i}k_{j}-\frac{1}{3}k^{2}\delta_{ij}\right)\delta_{m}(k),$ (26)
    where $\bar{C}_{1}=5\times 10^{-14}\mathrm{M}^{-1}_{\odot}h^{-2}\mathrm{Mpc}^{3}$ is a normalization constant, $\rho_{\mathrm{c}}=3H_{0}^{2}/8\pi G$ is the critical density, $D(z)$ is the growth factor, $z_{0}=0.62$ is a pivot redshift and $a_{1}$ and $\eta_{1}$ are model parameters.
@@
Cutoff $z\right>$ $0.33$ $0.54$ $0.74$ $1.01$ $1.62$
Cutoff 1 1.4 1.1 0.9 0.9 0.8
Cutoff 2 2.9 2.2 1.9 1.7 1.6
Cutoff 3 5.7 4.3 3.8 3.4 3.3
Table 6: Approximate scales in wavenumber $k$, measured in $h$Mpc-1, for each galaxy source bin that correspond to the angular cutoffs in $\xi_{+}$ tested in our LSST-like cosmic shear analysis. The scales are approximated by computing the angular diameter distance to the mean redshift of each bin and converting to a wavenumber.